# Batch-of-Thought: Cross-Instance Learning for Enhanced LLM Reasoning

### v1

@@ Abstract
Current Large Language Model reasoning systems process queries independently, discarding valuable cross-instance signals such as shared reasoning patterns and consistency constraints. We introduce Batch-of-Thought (BoT), a training-free method that processes related queries jointly to enable cross-instance learning. By performing comparative analysis across batches, BoT identifies high-quality reasoning templates, detects errors through consistency checks, and amortizes computational costs. We instantiate BoT within a multi-agent reflection architecture (BoT-R), where a Reflector performs joint evaluation to unlock mutual information gain unavailable in isolated processing. Experiments across three model families and six benchmarks demonstrate that BoT-R consistently improves accuracy and confidence calibration while reducing inference costs by up to 61%. Our theoretical and experimental analysis reveals when and why batch-aware reasoning benefits LLM systems.
@@ 1 Introduction
Large Language Models (LLMs) openai2023gpt4; madaan2023selfrefine; wei2022cot; shinn2023reflexion; yao2023react have achieved strong performance across diverse tasks and are increasingly applied in domains such as medical reasoning, question answering, and scientific problem solving (Singhal2025TowardEMA; mcduff2025towards; nori2025sequential; haas2025simpleqa; wang2023scibench; sun2024scieval). However, producing reliable answers with well-calibrated confidence remains a challenge  xiong2023uncertainty; ji2023hallucination_survey. LLMs often assign high confidence to incorrect answers, which undermines their practical deployment in high-stakes applications where accuracy and reliable uncertainty quantification are essential.
@@
Multi-agent LLM systems (li2023camel; chan2023chateval; guo2024large) extend single-model capabilities through specialized roles and iterative refinement. Despite architectural diversity, existing approaches share a fundamental limitation: they process queries independently. While computationally straightforward, this paradigm discards valuable cross-instance signals. When queries share domain characteristics or structural patterns, isolated processing prevents the system from identifying outliers through comparative assessment, propagating validated knowledge from confident instances to uncertain ones, or detecting errors that emerge only through cross-instance consistency checks.
@@
To illustrate this principle, consider fraud detection: evaluating a single seller in isolation provides limited signal. Examining a cohort simultaneously reveals recurring suspicious patterns, distinguishes legitimate domain practices from genuine anomalies, and enables comparative evidence assessment. This principle mirrors James–Stein estimation james1961quadratic; stein1956inadmissibility; efron1977steinsparadox: pooling information across similar instances improves individual estimates through shrinkage toward the cohort distribution (see Appendix C for theoretical analysis).
We instantiate BoT within a multi-agent reflection architecture madaan2023selfrefine; shinn2023reflexion, termed BoT-R, though the principle generalizes to other frameworks including Plan-and-Act Erdogan2025PlanAndAct and Multi-Agent Debate Du2023MultiAgentDebate; Liang2023MAD. In BoT-R, an Actor generates answer-rationale pairs for a batch of queries, then a Reflector performs joint evaluation through comparative analysis—identifying inconsistencies, extracting shared domain knowledge, assessing relative quality, and suggesting refinements. This approach simultaneously improves reasoning quality and computational efficiency by amortizing reflection overhead across the batch. We summarize our main contributions as follows:
@@ 2.1 Problem Formulation
A ReAct agent yao2023react that interleaves reasoning traces with tool execution to generate answer-rationale pairs. At iteration $t$, given query $x_{i}$ and optional critique $c_{i}^{(t-1)}$ from the previous round:
@@
A reflection agent madaan2023selfrefine; shinn2023reflexion that evaluates context $\mathcal{C}^{(t)}$ containing all current answer-rationale pairs. For each query $i$, it produces:
@@ 3.1 Experimental Setup
Our evaluation covers diverse reasoning and decision-making tasks: GPQA (rein2024gpqa), WinoGrande–debiased (sakaguchi2021winogrande), PubMedQA (jin2019pubmedqa), MedQA (USMLE) (jin2021disease), MMLU (hendrycks2020measuring), SMS Spam Detection (yang2024ad), and a newly curated dataset fraud-seller detection dataset (Appendix G). Together, these benchmarks span scientific reasoning, commonsense inference, biomedical QA, broad academic knowledge, and real-world anomaly detection.
@@
We evaluate (i) task accuracy, (ii) token efficiency (input token count, output token count, and total), and (iii) confidence calibration using two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation between the confidence distributions of correct vs. incorrect predictions, and (b) Expected Calibration Error (ECE; guo2017calibration).
@@
To isolate BoT’s contribution, we compare against two training-free reasoning baselines: ReAct (yao2023react), which performs standard single-instance reasoning with optional tool augmentation, and Reflection (shinn2023reflexion; madaan2023selfrefine), a multi-agent framework employing per-instance self-critique and revision. We then augment the same Actor with BoT’s joint, batch-aware reflection to obtain BoT-R. All other factors remain constant across methods.
@@
We report results from both API and open-source models: GPT-4o-2024-11-20 hurst2024gpt, Llama-3.3-70B llama3 and Qwen3-Next-80B yang2025qwen3. For the Fraud Detection dataset, we enable Brave Search and the Brave Summarizer as external tools for retrieval and grounding brave_api_docs
@@ Appendix A Experiment Settings
We use two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic smirnov1939estimation, which measures the maximum difference between the cumulative distributions of confidence scores for correct versus incorrect predictions—higher KS values indicate better separation and thus more reliable confidence estimates; and (b) Expected Calibration Error (ECE guo2017calibration), which quantifies the average gap between predicted confidence and actual accuracy across binned predictions—lower ECE indicates that confidence scores accurately reflect true correctness probabilities. Together, these metrics assess both discriminative power (KS) and absolute calibration quality (ECE).
@@ D.1 Confidence Calibration in Large Language Models
Reliable uncertainty quantification remains critical for deploying LLMs in high-stakes applications. Modern LLMs frequently exhibit poor calibration, assigning high confidence to incorrect predictions (guo2017calibration; xiong2023uncertainty; kadavath2022language). This miscalibration persists even in state-of-the-art models (openai2023gpt4; anthropic2024claude), undermining trust in automated decision-making systems.
Existing calibration approaches fall into three categories. Post-hoc calibration methods apply temperature scaling (guo2017calibration) or Platt scaling to model outputs, but require held-out calibration sets and fail to capture semantic uncertainty (kuhn2023semantic). Sampling-based methods estimate uncertainty through self-consistency (wang2023selfconsistency), semantic entropy (kuhn2023semantic), or ensemble disagreement (chen2023reconcile). While effective, these approaches incur substantial computational overhead—self-consistency requires 10-40 samples per query—and process each instance independently, missing opportunities for cross-instance calibration. Verbalized confidence approaches directly prompt models for numerical (xiong2023uncertainty; lin2022teaching) or categorical (tian2023justsay) confidence estimates. These methods are efficient but highly sensitive to prompt formatting (si2024prompting) and often produce overconfident predictions (kadavath2022language). Recent efforts employ chain-of-thought reasoning for confidence elicitation (xiong2023uncertainty) or fine-tune models on calibration data (lin2022teaching), yet these remain per-instance techniques that cannot leverage distributional signals.
@@ D.2 Multi-Agent Reasoning Systems
Recent work has explored sophisticated communication protocols (wu2023autogen), dynamic role allocation (hong2023metagpt), and multi-agent collaboration on complex tasks (li2023camel; qian2024chatdev). However, a fundamental limitation persists: existing multi-agent systems process queries independently. Even when multiple agents collaborate on a single query, the framework treats each query in isolation, discarding cross-instance signals. AutoGen (wu2023autogen) and MetaGPT (hong2023metagpt) enable multi-agent workflows but apply them instance-by-instance. CAMEL (li2023camel) studies role-playing conversations yet maintains per-query boundaries.
The closest work to ours is batch prompting (cheng2023batch), which groups multiple queries into a single API call for efficiency. However, batch prompting lacks reflective evaluation mechanisms and does not perform comparative analysis—it simply concatenates queries without leveraging cross-instance reasoning. Our work fundamentally differs by introducing batch-aware reflection: the Reflector explicitly performs comparative evaluation, consistency checking, and knowledge propagation across the batch.
@@ D.3 Cross-Instance Learning
In deep learning, batch normalization (ioffe2015batch) leverages mini-batch statistics during training, while recent work explores cross-example attention (lee2019set) and set-based reasoning (zaheer2017deep). Meta-learning approaches (finn2017maml; snell2017prototypical) learn from task distributions rather than individual instances, demonstrating benefits of comparative learning.
For LLM inference, in-context learning (brown2020language) uses examples to guide reasoning, and analogical prompting (yasunaga2024large) retrieves similar cases to aid problem-solving. However, these methods rely on predefined examples or retrieval systems rather than jointly reasoning over a batch of target queries. Recent work on self-consistency with rationalization (mitchell2023enhancing) aggregates multiple reasoning paths for a single query but does not transfer knowledge across distinct queries.
@@ D.4 Positioning of Our Work
(3) Training-free generality: Unlike calibration methods requiring fine-tuning (lin2022teaching) or specialized architectures, BoT is model-agnostic and integrates with existing multi-agent frameworks (Reflection, Plan-and-Act, Debate) without additional training.
@@ E.1 Batch Size Effects
We further study how batch composition influences BoT-R by comparing three batching strategies across six datasets. No-batch applies reflection independently to each query without any cross-instance context. Sequential batching groups queries in their original dataset order—without explicitly enforcing semantic similarity—and is used as the default strategy in our main experiments. Semantic batching clusters queries by embedding similarity using K-means over E5-Mistral-7B embeddings (wang2023improving), and forms fixed-size batches from cluster members sorted by proximity to the cluster centroid to maximize within-batch coherence. Results are summarized in Table 7.

### v2

@@ Abstract
Current Large Language Model reasoning systems process queries independently, discarding valuable cross-instance signals such as shared reasoning patterns and consistency constraints. We introduce Batch-of-Thought (BoT), a training-free method that processes related queries jointly to enable cross-instance learning. By performing comparative analysis across batches, BoT identifies high-quality reasoning templates, detects errors through consistency checks, and amortizes computational costs. We instantiate BoT within a multi-agent reflection architecture (BoT-R), where a Reflector performs joint evaluation to unlock mutual information gain unavailable in isolated processing. Experiments across three model families and six benchmarks demonstrate that BoT-R consistently improves accuracy and confidence calibration while reducing inference costs by up to 61%. Our theoretical and experimental analysis reveals when and why batch-aware reasoning benefits LLM systems. Our code is available at https://github.com/xuanyang19/BoT.
@@ 1 Introduction
Large Language Models (LLMs) OpenAI (2023); Madaan et al. (2023); Wei et al. (2022); Shinn et al. (2023); Yao et al. (2023) have achieved strong performance across diverse tasks and are increasingly applied in domains such as medical reasoning, question answering, and scientific problem solving (Singhal et al., 2025; McDuff et al., 2025; Nori et al., 2025; Haas et al., 2025; Wang et al., 2023b; Sun et al., 2024). However, producing reliable answers with well-calibrated confidence remains a challenge  Xiong et al. (2023); Ji et al. (2023). LLMs often assign high confidence to incorrect answers, which undermines their practical deployment in high-stakes applications where accuracy and reliable uncertainty quantification are essential.
@@
Multi-agent LLM systems (Li et al., 2023; Chan et al., 2023; Guo et al., 2024) extend single-model capabilities through specialized roles and iterative refinement. Despite architectural diversity, existing approaches share a fundamental limitation: they process queries independently. While computationally straightforward, this paradigm discards valuable cross-instance signals. When queries share domain characteristics or structural patterns, isolated processing prevents the system from identifying outliers through comparative assessment, propagating validated knowledge from confident instances to uncertain ones, or detecting errors that emerge only through cross-instance consistency checks.
@@
To illustrate this principle, consider fraud detection: evaluating a single seller in isolation provides limited signal. Examining a cohort simultaneously reveals recurring suspicious patterns, distinguishes legitimate domain practices from genuine anomalies, and enables comparative evidence assessment. This principle mirrors James–Stein estimation James and Stein (1961); Stein (1956); Efron and Morris (1977): pooling information across similar instances improves individual estimates through shrinkage toward the cohort distribution (see Appendix C for theoretical analysis).
We instantiate BoT within a multi-agent reflection architecture Madaan et al. (2023); Shinn et al. (2023), termed BoT-R, though the principle generalizes to other frameworks including Plan-and-Act Erdogan et al. (2025) and Multi-Agent Debate Du et al. (2023); Liang et al. (2023). In BoT-R, an Actor generates answer-rationale pairs for a batch of queries, then a Reflector performs joint evaluation through comparative analysis—identifying inconsistencies, extracting shared domain knowledge, assessing relative quality, and suggesting refinements. This approach simultaneously improves reasoning quality and computational efficiency by amortizing reflection overhead across the batch. We summarize our main contributions as follows:
@@ 2.1 Problem Formulation
A ReAct agent Yao et al. (2023) that interleaves reasoning traces with tool execution to generate answer-rationale pairs. At iteration $t$, given query $x_{i}$ and optional critique $c_{i}^{(t-1)}$ from the previous round:
@@
A reflection agent Madaan et al. (2023); Shinn et al. (2023) that evaluates context $\mathcal{C}^{(t)}$ containing all current answer-rationale pairs. For each query $i$, it produces:
@@ 3.1 Experimental Setup
Our evaluation covers diverse reasoning and decision-making tasks: GPQA (Rein et al., 2024), WinoGrande–debiased (Sakaguchi et al., 2021), PubMedQA (Jin et al., 2019), MedQA (USMLE) (Jin et al., 2021), MMLU (Hendrycks et al., 2020), SMS Spam Detection (Yang et al., 2024), and a newly curated dataset fraud-seller detection dataset (Appendix G). Together, these benchmarks span scientific reasoning, commonsense inference, biomedical QA, broad academic knowledge, and real-world anomaly detection.
@@
We evaluate (i) task accuracy, (ii) token efficiency (input token count, output token count, and total), and (iii) confidence calibration using two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic Smirnov (1939) between the confidence distributions of correct vs. incorrect predictions, and (b) Expected Calibration Error (ECE; Guo et al. (2017)).
@@
To isolate BoT’s contribution, we compare against two training-free reasoning baselines: ReAct (Yao et al., 2023), which performs standard single-instance reasoning with optional tool augmentation, and Reflection (Shinn et al., 2023; Madaan et al., 2023), a multi-agent framework employing per-instance self-critique and revision. We then augment the same Actor with BoT’s joint, batch-aware reflection to obtain BoT-R. All other factors remain constant across methods.
@@
We report results from both API and open-source models: GPT-4o-2024-11-20 Hurst et al. (2024), Llama-3.3-70B Dubey et al. (2024) and Qwen3-Next-80B Yang et al. (2025). For the Fraud Detection dataset, we enable Brave Search and the Brave Summarizer as external tools for retrieval and grounding Brave Software, Inc. (2025)
@@ Appendix A Experiment Settings
We use two complementary measures: (a) the Kolmogorov–Smirnov (KS) statistic Smirnov (1939), which measures the maximum difference between the cumulative distributions of confidence scores for correct versus incorrect predictions—higher KS values indicate better separation and thus more reliable confidence estimates; and (b) Expected Calibration Error (ECE Guo et al. (2017)), which quantifies the average gap between predicted confidence and actual accuracy across binned predictions—lower ECE indicates that confidence scores accurately reflect true correctness probabilities. Together, these metrics assess both discriminative power (KS) and absolute calibration quality (ECE).
@@ D.1 Confidence Calibration in Large Language Models
Reliable uncertainty quantification remains critical for deploying LLMs in high-stakes applications. Modern LLMs frequently exhibit poor calibration, assigning high confidence to incorrect predictions (Guo et al., 2017; Xiong et al., 2023; Kadavath et al., 2022). This miscalibration persists even in state-of-the-art models (OpenAI, 2023; Anthropic, 2024), undermining trust in automated decision-making systems.
Existing calibration approaches fall into three categories. Post-hoc calibration methods apply temperature scaling (Guo et al., 2017) or Platt scaling to model outputs, but require held-out calibration sets and fail to capture semantic uncertainty (Kuhn et al., 2023). Sampling-based methods estimate uncertainty through self-consistency (Wang et al., 2023c), semantic entropy (Kuhn et al., 2023), or ensemble disagreement (Chen et al., 2024). While effective, these approaches incur substantial computational overhead—self-consistency requires 10-40 samples per query—and process each instance independently, missing opportunities for cross-instance calibration. Verbalized confidence approaches directly prompt models for numerical (Xiong et al., 2023; Lin et al., 2022) or categorical (Tian et al., 2023) confidence estimates. These methods are efficient but highly sensitive to prompt formatting (Si et al., 2024) and often produce overconfident predictions (Kadavath et al., 2022). Recent efforts employ chain-of-thought reasoning for confidence elicitation (Xiong et al., 2023) or fine-tune models on calibration data (Lin et al., 2022), yet these remain per-instance techniques that cannot leverage distributional signals.
@@ D.2 Multi-Agent Reasoning Systems
Recent work has explored sophisticated communication protocols (Wu et al., 2023), dynamic role allocation (Hong et al., 2023), and multi-agent collaboration on complex tasks (Li et al., 2023; Qian et al., 2024). However, a fundamental limitation persists: existing multi-agent systems process queries independently. Even when multiple agents collaborate on a single query, the framework treats each query in isolation, discarding cross-instance signals. AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2023) enable multi-agent workflows but apply them instance-by-instance. CAMEL (Li et al., 2023) studies role-playing conversations yet maintains per-query boundaries.
The closest work to ours is batch prompting (Cheng et al., 2023), which groups multiple queries into a single API call for efficiency. However, batch prompting lacks reflective evaluation mechanisms and does not perform comparative analysis—it simply concatenates queries without leveraging cross-instance reasoning. Our work fundamentally differs by introducing batch-aware reflection: the Reflector explicitly performs comparative evaluation, consistency checking, and knowledge propagation across the batch.
@@ D.3 Cross-Instance Learning
In deep learning, batch normalization (Ioffe and Szegedy, 2015) leverages mini-batch statistics during training, while recent work explores cross-example attention (Lee et al., 2019) and set-based reasoning (Zaheer et al., 2017). Meta-learning approaches (Finn et al., 2017; Snell et al., 2017) learn from task distributions rather than individual instances, demonstrating benefits of comparative learning.
For LLM inference, in-context learning (Brown et al., 2020) uses examples to guide reasoning, and analogical prompting (Yasunaga et al., 2024) retrieves similar cases to aid problem-solving. However, these methods rely on predefined examples or retrieval systems rather than jointly reasoning over a batch of target queries. Recent work on self-consistency with rationalization (Mitchell et al., 2023) aggregates multiple reasoning paths for a single query but does not transfer knowledge across distinct queries.
@@ D.4 Positioning of Our Work
(3) Training-free generality: Unlike calibration methods requiring fine-tuning (Lin et al., 2022) or specialized architectures, BoT is model-agnostic and integrates with existing multi-agent frameworks (Reflection, Plan-and-Act, Debate) without additional training.
@@ E.1 Batch Size Effects
We further study how batch composition influences BoT-R by comparing three batching strategies across six datasets. No-batch applies reflection independently to each query without any cross-instance context. Sequential batching groups queries in their original dataset order—without explicitly enforcing semantic similarity—and is used as the default strategy in our main experiments. Semantic batching clusters queries by embedding similarity using K-means over E5-Mistral-7B embeddings (Wang et al., 2023a), and forms fixed-size batches from cluster members sorted by proximity to the cluster centroid to maximize within-batch coherence. Results are summarized in Table 7.